OpenStudy (anonymous):

simplify the expression "the square root of x^10 times the square root of x^4." a. the square root of x^14 b. 2 times the square root of x^10 d. x^7

OpenStudy (anonymous):

(x^10)^0.5*(x^4)^0.5 =x^(10*0.5)*x^(4*0.5) =x^5*x^2 =x^(5+2) =x^7

OpenStudy (anonymous):

so what's the answer?

OpenStudy (anonymous):

d

OpenStudy (anonymous):

sorry! i couldn't understand it